A body travels 200cm in first  2s and 220cm in the next 4s . What will be the velocity at the end of 7th second from start?

Grade:11

1 Answers

Soumya Ranjan Mohanty
6 years ago
Dear student,
see given displacement in 2 sec=200 cm
let initial velocity=u
putting s=ut + 1/2 at^2
=>200=u*2 + 05 a*2^2
=>100=a+u--(1)
now for t=6, equation becomes,
420=u*6+0.5a*6^2
=>70=3a+u --(2)
solving (1) and (2),
a=-15 cm/sec^2
putting a value in (1) or (2) we get, u= 115 cm/sec
now when t=7sec, we use v=u+at
=>v=115-15*7= 10 cm/sec [Ans]
Regards,
Thank you,
Soumya Ranjan Mohanty.